### v1

@@ 1 Introduction
Our goal is to contribute to the extensive research developed over the past decades by implementing and testing a special class of wavelet transforms based on a non-standard multiresolution analysis in which the resolution level is tripled instead of doubled at each step. Specifically, we consider a parametric family of wavelets recently introduced in [12], constructed without any dilation and translation but using de la Vallée Poussin (briefly VP) interpolation at Chebyshev nodes [13, 14, 15]. These VP wavelets are interpolating, non-orthogonal, polynomials. They are defined on $[-1,1]$, to which any other compact interval can be mapped, and can be easily generalized to higher dimensions via tensor products. Thus, unlike classical wavelets, data on finite domains can be naturally processed by VP wavelets without the need for additional modifications. Furthermore, even with the uniform norm, they possess good approximation properties, which can be further controlled through a free parameter.
@@ 2 Univariate VP scaling and wavelet functions
$\phi_{n}^{m}(x,y)=\frac{2}{n}\sum_{r=0}^{n+m-1}{}^{\prime}\mu_{n,r}^{m}T_{r}(x)T_{r}(y),\qquad n,m\in\mathbb{N},\quad 0<m<n,$ (1)
@@
[…]
@@
    Convergence. For any fixed $\theta\in]0,1[$, if we take $m=\lfloor\theta n\rfloor$ and a sufficiently large resolution level $n\in\mathbb{N}$, then any continuous function $f$ on $[-1,1]$ can be uniformly approximated with the desired precision by the VP interpolation polynomial [14]
@@
    Furthermore, if $f$ is continuously differentiable up to order $2s$, for some $s\in\mathbb{N}$, then we have [11, Thm. 5.3]
@@ 3 Decomposition and reconstruction algorithms
The above defined VP scaling and wavelet functions generate a multiresolution analysis, whose approximation and detail spaces, at a given resolution level $n\in\mathbb{N}$ and arbitrarily fixed parameter $m<n$, are defined as follows
$\mathscr{V}_{n}^{m}$ $=$ $\mbox{span}\,\{\phi_{n,k}^{m},\ k=1,\ldots,n\},\qquad\dim\mathscr{V}_{n}^{m}=n,$ (14)
$\mathscr{W}_{n}^{m}$ $=$ $\mbox{span}\,\{\psi_{n,k}^{m},\ k=1,\ldots,2n\},\qquad\dim\mathscr{W}_{n}^{m}=2n.$ (15)
They satisfy [12]
$\mathscr{V}^{m}_{3n}=\mathscr{V}^{m}_{n}\oplus\mathscr{W}_{n}^{m}\quad\mbox{and}\quad\mathscr{V}_{n}^{m}\bot\mathscr{W}_{n}^{m},\qquad\quad\forall n>m\in\mathbb{N},$ (16)
where orthogonality is understood w.r.t. the scalar product
$<f,g>:=\int_{-1}^{1}f(x)g(x)dw(x),\qquad dw(x)=\frac{dx}{\sqrt{1-x^{2}}}.$
Consequently, any function $f_{3n}\in\mathscr{V}_{3n}^{m}$ can be uniquely decomposed as
@@
Fast decomposition and reconstruction algorithms can be given in order to compute the basis coefficients of $f_{n}$ and $g_{2n}$ from the basis coefficients of $f_{3n}$, and viceversa [12, Thm. 5.2].
@@
6:  $\vec{x}=\frac{\pi}{3n}\mathscr{T}_{1}\left(\vec{v}\right)$
7:  $\vec{y}=\frac{\pi}{3n}\mathscr{T}_{3}\left(\vec{u}\right)$
8:  for $k=1:n-m+1$ do
@@
14:  $\vec{a}_{n}=\mathscr{T}_{2}(\vec{v})$
15:  $\vec{b}_{n}=\mathscr{T}_{4}(\vec{v})$
Algorithm 1 Decomposition
@@
1:  $\vec{x}=\frac{\pi}{n}\mathscr{T}_{1}\left(\vec{a}_{n}\right)$
2:  $\vec{y}=\frac{\pi}{3n}\mathscr{T}_{3}\left(\vec{b}_{n}\right)$
3:  $\vec{u}=\vec{a}_{n}-\mathscr{T}_{2}(\vec{y})$
4:  $\vec{v}=\vec{b}_{n}+\mathscr{T}_{4}(\vec{x})$
@@
Algorithm 2 Reconstruction
As observed in [12], each transform $\mathscr{T}_{i}$ can be efficiently computed using $\mathscr{O}(n\log n)$ flops by means of the well–known discrete cosine transform (DCT) and its inverse (IDCT), in Matlab known as DCT of type 2 and 3, respectively. They transform any $\vec{v}=(v_{1},\ldots,v_{N})\in{\mathbb{R}}^{N}$ in $\vec{u}\in{\mathbb{R}}^{N}$ defined as follows
@@
Moreover, the transforms $\mathscr{T}_{3}$ and $\mathscr{T}_{4}$ can be computed using the following algorithms. Input: $\vec{v}=(v_{1},\ldots,v_{2n})$ Output: $\vec{u}=(u_{1},\ldots,u_{n})$ 1:  for $k=1:n$ do 2:     $w_{3k-2}=v_{2k-1}$ 3:     $w_{3k-1}=0$ 4:     $w_{3k}=v_{2k}$ 5:  end for 6:  $\vec{x}=\sqrt{\frac{3n}{\pi}}DCT(\vec{w})$ 7:  for $k=1:n-m+1$ do 8:     $u_{k}=x_{k}$ 9:  end for 10:  for $k=n-m+2:n$ do 11:     $u_{k}=\frac{m+(n-k+1)}{2m}x_{k}-\frac{m-(n-k+1)}{2m}x_{2n-k+2}$ 12:  end for Algorithm 3 Compute $\vec{u}=\mathscr{T}_{3}(\vec{v})$    Input: $\vec{v}=(v_{1},\ldots,v_{n})$ Output: $\vec{u}=(u_{1},\ldots,u_{2n})$ 1:  for $k=1:n-m+1$ do 2:     $w_{k}=v_{k}$ 3:  end for 4:  for $k=n-m+2:n$ do 5:     $w_{k}=\frac{m+(n-k+1)}{2m}v_{k}$ 6:     $w_{2n-k+2}=-\frac{m-(n-k+1)}{2m}v_{k}$ 7:  end for 8:  for $k=n+m+4:3n$ do 9:     $w_{k}=0$ 10:  end for 11:  $\vec{x}=\sqrt{\frac{3n}{\pi}}IDCT(\vec{w})$ 12:  for $k=1:n$ do 13:     $u_{2k-1}=x_{3k-2}$ 14:     $u_{2k}=x_{3k}$ 15:  end for Algorithm 4 Compute $\vec{u}=\mathscr{T}_{4}(\vec{v})$
3.1 Application to signal processing
The application of VP bases to signal processing requires some considerations. First of all, let us discuss about the mathematical model to adopt. As is well known, a digital signal is mathematically represented by a vector, say $\vec{v}=(v_{1},\ldots,v_{N})$, of a certain length $N$. Its elements are typically considered as the values assumed by a certain function on $N$ equidistant nodes of an interval, which we can assume to be $[-1,1]$ without violating generality. In our model, we consider equidistant nodes not on the segment $[-1,1]$ but on the semicircle, i.e.
@@ 3.1 Application to signal processing
$v_{k}=f(\cos t_{k}),\qquad k=1,\ldots,N.$
Thus, using these samples we can approximate $f$, and hence the signal, by its projection onto the space $\mathscr{V}_{N}$, i.e., by the VP interpolation polynomial
$V_{N}^{m}f(x)=\sum_{k=1}^{N}v_{k}\phi_{N,k}^{m}(x),\quad m=\lfloor\theta N\rfloor,\quad\theta\in]0,1[,\qquad x\in[-1,1],$
and we can handle the signal with the usual wavelet techniques by applying the previous decomposition and reconstruction algorithms.
We now have to face two problems: the first concerns $N$, which according to our scheme should be a multiple of three, the second concerns the conservation of the signal energy.
To address the first issue, we extend the vector $\vec{v}$ by replicating its last element until its length becomes divisible by three, after which we apply the decomposition algorithm. During reconstruction, the vector of scaling coefficients is then truncated to restore the original length.
Regarding the second problem, since the VP scaling and wavelet functions are interpolating and not orthogonal, the Parseval identity is not applicable, and appropriate correction factors must be considered both in decomposition and reconstruction for energy conservation. Traditionally, in case of non-orthogonal wavelets, the scaling and wavelet coefficients are usually multiplied by $\sqrt{2}$ after the decomposition, and divided by $\sqrt{2}$ before the reconstruction. In our case, we multiply/divide the scaling coefficients by $\sqrt{3}$ and the wavelet coefficients by $\sqrt{3/2}$. To justify this, we observe that $f_{3n}=f_{n}+g_{2n}$ and $f_{n}\bot g_{2n}$ imply
@@
Figure 1: Six wavelet test functions: Blocks, Bumps, Heavy Sine, Doppler, Quadchirp, Mishmash.
Figure 2: The results for different values of $\theta=0.1,0.2,\ldots,0.9$ for the six wavelet test functions: Blocks, Bumps, Heavy Sine, Doppler, Quadchirp, Mishmash.
@@
With the described algorithms and all the above observations in mind, we now proceed to test the VP basis through experiments on two typical data processing tasks: signal denoising and image compression. In the first case, we use univariate VP wavelets, in the second case we deal with bidimensional VP wavelets. The bivariate extension of the VP wavelets is obtained through a standard tensor product approach of the univariate scaling and wavelet functions. According to the previous semicircular sampling model, it is assumed that the equidistant pixel grid of the image is mapped to the grid of Chebyshev nodes. Following the standard procedure, the algorithms described above are applied iteratively to the columns and rows of the image data.

### v2

@@ 1 Introduction
Our goal is to contribute to the extensive research developed over the past decades by implementing and testing a special class of wavelet transforms based on a non-standard multiresolution analysis in which the resolution level is tripled instead of doubled at each step. Specifically, we consider a parametric family of polynomial wavelets recently introduced in [19]. Similarly to other kinds of polynomial wavelets (see, e.g. [10, 11, 12]) there is no corresponding scaling or wavelet mother functions neither any dilation and translation is implemented, but, at any resolution level, we have different, well localized, polynomial scaling and wavelet basis functions generating the appproximation and detail spaces, respectively.
The polynomial wavelets examined in this paper are based on de la Vallée Poussin (VP) interpolation at Chebyshev nodes [20, 21, 22]. These VP wavelets are interpolating, non-orthogonal polynomials defined on the interval [1,1][-1,1], to which any other compact interval can be mapped. They can also be naturally extended to higher dimensions through tensor-product constructions. Hence, data defined on finite domains can be processed directly by VP wavelets, without requiring additional modifications, unlike classical wavelets, which are intrinsically defined on the entire real line through translations and dilations of a single mother function..
Furthermore, even with the uniform norm, they possess good approximation properties, which can be further controlled through a free parameter.
@@ 2 Univariate VP scaling and wavelet functions
$\phi_{n}^{m}(x,y)=\frac{2}{n}\sum_{r=0}^{n+m-1}\hskip-2.84544pt{}^{\prime}\mu_{n,r}^{m}T_{r}(x)T_{r}(y),\qquad n,m\in\mathbb{N},\quad 0<m<n,$ (1)
@@
Note that the terms “scaling function” and “wavelet” are used here in a broader sense. Although these functions are not generated by dyadic dilations and translations of a single prototype function, they share several key wavelet-like features, as illustrated below. This is in line with earlier polynomial constructions already called ”wavelets” (e.g., [10, 11, 12]).
@@
    $\|f\|_{p}=\left\{\begin{array}[]{ll}\displaystyle\left(\int_{-1}^{1}|f(x)|^{p}w(x)dx\right)^{\frac{1}{p}}&1\leq p<\infty,\\[7.22743pt] \displaystyle\sup_{|x|\leq 1}|f(x)|&p=\infty,\end{array}\right$.
@@
    Convergence. For any fixed $\theta\in]0,1[$, if we take $m=\lfloor\theta n\rfloor$ and a sufficiently large resolution level $n\in\mathbb{N}$, then any continuous function $f$ on [1,1][-1,1] can be uniformly approximated with the desired precision by the VP interpolation polynomial [21]
@@
    Furthermore, if $f$ is continuously differentiable up to order $2s$, for some $s\in\mathbb{N}$, then we have [18, Thm. 5.3]
@@ 3 Decomposition and reconstruction algorithms
The above defined VP scaling and wavelet functions generate a hierarchical structure of approximation and detail spaces. At a given resolution level $n\in\mathbb{N}$ and for an arbitrarily fixed parameter $m<n$, we define
$\mathscr{V}_{n}^{m}$ $=$ $\operatorname{span}\,\{\phi_{n,k}^{m},\ k=1,\ldots,n\},\qquad\dim\mathscr{V}_{n}^{m}=n,$ (14)
$\mathscr{W}_{n}^{m}$ $=$ $\operatorname{span}\,\{\psi_{n,k}^{m},\ k=1,\ldots,2n\},\qquad\dim\mathscr{W}_{n}^{m}=2n.$ (15)
Each space $\mathscr{V}_{n}^{m}$ is finite-dimensional and therefore a closed subspace of the weighted Hilbert space $L^{2}_{w}([-1,1])$, endowed with the scalar product
$\langle f,g\rangle:=\int_{-1}^{1}f(x)g(x)\,dw(x),\qquad dw(x)=\frac{dx}{\sqrt{1-x^{2}}}.$
Moreover, these spaces satisfy [19]
$\mathscr{V}^{m}_{3n}=\mathscr{V}^{m}_{n}\oplus\mathscr{W}_{n}^{m},\qquad\mathscr{V}_{n}^{m}\perp\mathscr{W}_{n}^{m},\qquad\forall\,n>m\in\mathbb{N}.$ (16)
Remark. The relation (16) defines a nested sequence of polynomial approximation and detail spaces that mirrors the hierarchical refinement of a multiresolution analysis (MRA). However, since no translation or dilation operators are involved and $\mathscr{V}_{n}^{m}$ is not a shift-invariant subspace of $L^{2}(\mathbb{R})$, this construction does not yield a ”classical” multiresolution analysis (MRA). In this paper, this term is therefore used in a nonstandard or generalized sense to highlight the hierarchical and orthogonal decomposition (16) which refines the resolution by a factor of three on the compact interval [1,1][-1,1].
In virtue of (16), any function $f_{3n}\in\mathscr{V}_{3n}^{m}$ can be uniquely decomposed as
@@
Fast decomposition and reconstruction algorithms can be given in order to compute the basis coefficients of $f_{n}$ and $g_{2n}$ from the basis coefficients of $f_{3n}$, and viceversa [19, Thm. 5.2].
@@
.2cm
[…]
.1cm
[…]
.2cm
8:for k=1:nm+1k=1:n-m+1 do
.1cm
@@
[…]
Algorithm 1 Decomposition (classic)
@@
.2cm
[…]
.2cm
[…]
.2cm
[…]
.2cm
[…]
.2cm
@@
Algorithm 2 Reconstruction (classic)
As observed in [19], each transform $\mathscr{T}_{i}$ can be efficiently computed using the well–known discrete cosine transform (DCT) and its inverse (IDCT), in Matlab known as DCT of type 2 and 3, respectively. They transform any $\vec{v}=(v_{1},\ldots,v_{n})\in{\mathbb{R}}^{n}$ in $\vec{u}\in{\mathbb{R}}^{n}$ defined as follows
@@
[…]
We conclude the section with some remarks about the computational complexity. The proposed transform requires $O(N\log N)$ operations per level due to its DCT-based implementation, compared with $O(N)$ for standard dyadic wavelets. However, its ternary decomposition structure ($\mathscr{V}_{3n}^{m}=\mathscr{V}_{n}^{m}\oplus\mathscr{W}_{n}^{m}$) allows more substantial data reduction per step, resulting in fewer decomposition levels and an overall computational cost that remains practically competitive while maintaining reconstruction quality, as confirmed in the following experimental sections.
4 Application to signal processing
The application of VP bases to signal processing requires some considerations.
First of all, let us discuss the mathematical model to adopt. As it is well known, a digital signal is mathematically represented by a vector, say $\vec{v}=(v_{1},\ldots,v_{N})$, of a certain length $N$. Its elements are typically considered as the values assumed by a certain function on $N$ equidistant nodes of an interval, which we can assume to be [1,1][-1,1] without violating generality. In our model, we consider equidistant nodes not on the segment [1,1][-1,1] but on the semicircle, i.e.
@@ 4 Application to signal processing
$v_{k}=f(\cos t_{k}),\qquad k=1,\ldots,N.$ (20)
Thus, using these samples we can approximate $f$, and hence the signal, by its VP interpolation polynomial
$V_{N}^{m}f(x)=\sum_{k=1}^{N}v_{k}\phi_{N,k}^{m}(x),\quad m=\lfloor\theta N\rfloor,\quad\theta\in]0,1[,\qquad x\in[-1,1],$ (21)
which represents the interpolating projection onto the approximation space $\mathscr{V}_{N}^{m}$, for arbitrary values of the parameter $\theta$, and hence $m$.
The non-uniform sampling (20) at the Chebyshev zeros and the consequent VP approximation (21) of a digital signal $\vec{v}=(v_{1},\ldots,v_{N})$ represent the non standard mathematical model adopted here. We recall that this model has been recently applied in [14, 15] where, taking the multivariate extension via tensor product, the authors get efficient and competitive algorithms for zooming arbitrary 2D and 3D images just resampling the multivariate VP approximation at the Chebyshev grid of the desired size.
In this paper, we start from the same sampling/approximation model with the aim of handling the signal $\vec{v}$, interpreted as a scaling vector at the resolution $N$, with the usual wavelet techniques: the decomposition and reconstruction algorithms, respectively described in Algorithm 1 and Algorithm 2.
To this aim, now, we have to face two problems: the first concerns $N$, which according to our scheme should be a multiple of three, the second concerns the conservation of the signal energy.
To address the first issue, before applying the decomposition step described in Algorithm 1, we enlarge the input scaling vector $\vec{v}$ whenever its length is not a multiple of three, by adding up to two extra elements so that its size becomes divisible by three.
Different strategies can be used to define these added elements. In Algorithm 5, we introduce a parameter ind controlling the extension rule: if ind=0=0, the added elements are set to 0; if ind=1=1, they replicate the last entry $v_{N}$ of $\vec{v}$. Since the VP approximation preserves constants, to avoid introducing spurious oscillations we fixed ind=1=1 in all the experiments reported in this paper. However, when we set ind=0=0, the numerical results were essentially unchanged. In fact, because at most two elements are added before each decomposition step, the effect of the chosen rule becomes negligible, especially for large input sizes.
During reconstruction, the previously added elements are removed to restore the original signal length. This truncation is performed level by level: before each reconstruction step, we check whether the length of the scaling vector is twice that of the corresponding detail vector. If not, we deduce that the scaling vector had been enlarged during decomposition, and we discard its last added entries so that the length ratio is restored. After the final reconstruction, the recovered signal may differ from the original by at most two samples, which corresponds to a very small information loss. In practice, the effect of this truncation on both the numerical and visual results was found to be insignificant.
Regarding the second problem, since the VP scaling and wavelet functions are interpolating and not orthogonal, the Parseval identity is not applicable, and appropriate correction factors must be considered both in decomposition and reconstruction in order to assure energy conservation. In the classical wavelet setting, the scaling and wavelet coefficients are usually multiplied by $\sqrt{2}$ in the decomposition step, and divided by $\sqrt{2}$ in the reconstruction step. In our case, we multiply/divide the scaling coefficients by $\sqrt{3}$ and the wavelet coefficients by $\sqrt{3/2}$. To justify this, we observe that $f_{3n}=f_{n}+g_{2n}$ and $f_{n}\bot g_{2n}$ imply
@@
With all the above observations in mind, in the following Algorithm 5 and Algorithm 6, we summarize all the adjustments to the decomposition and reconstruction Algorithms 1 and 2, respectively. As we have previously discussed, these algorithms have no assumption on the size, which can also be not a multiple of three. Moreover, we fix by default $ind=1$ regarding the method to enlarge the scaling vectors, and we use the normalization factors $f_{sca}=\sqrt{3}$ and $f_{wav}=\sqrt{3/2}$ to multiply (in decomposition) and divide (in reconstruction) with the scaling and detail vectors, respectively.
Input: $\vec{v}=(v_{1},\ldots,v_{N})$ scaling vector to decompose
Output: $\vec{u}$ scaling vector
      w\ $\vec${w} detail vector
.2cm
1:$ind=1$; $f_{sca}=\sqrt{3}$; $f_{wav}=\sqrt{\frac{3}{2}}$
2:$R=$ remainder term of $N/3$
3:if $R\neq 0$ then
4:  for k=1:3Rk=1:3-R do
5:   if $ind=1$ then
6:    $v_{N+k}=v_{N}$
7:   else
8:    $v_{N+k}=0$
9:   end if
10:  end for
11:end if
12: Run Algorithm 1 with input vector $\vec{a}_{3n}=\vec{v}$ obtaining as output the scaling vector $\vec{a}_{n}$ and the detail vector $\vec{b}_{2n}$
[…]
Algorithm 5 Decomposition (any size)
Input: $\vec{u}$ scaling vector
      w\ $\vec${w} detail vector
Output: $\vec{v}=(v_{1},\ldots,v_{N})$ reconstructed scaling vector
.2cm
[…]
$2n=length(\vec{w})/$2
[…]
6: Run Algorithm 2 with input vectors $\vec{a}_{n}$ and $\vec{b}_{2n}$ obtaining as output the reconstructed scaling vector $\vec{v}$
Algorithm 6 Reconstruction (any size)
In Figure 1 a comparison is shown between the choices $f_{sca}=\sqrt{3}$, $f_{wav}=\sqrt{\frac{3}{2}}$ and $f_{sca}=f_{wav}=1$. Note that for $f_{sca}=\sqrt{3}$, $f_{wav}=\sqrt{\frac{3}{2}}$ the sum of the squared norms is almost equal to $1$.
Figure 1: The mean and standard deviation of the sum of the squared norms for $100$ samples for a signal of length $3^{7}$, $\theta=0.5$ and 77 steps done for $f_{sca}=\sqrt{3}$, $f_{wav}=\sqrt{\frac{3}{2}}$ (left) and for $f_{sca}=f_{wav}=1$ (right).
Let us conclude the section with some considerations concerning the extension of the previous arguments to the bivariate case that is of interest in the case that the signal to process is an image.
As is well–known, mathematically, any digital (gray-level) image at the resolution $N\times M$ is a matrix $I$ with size $N\times M$. According to the mathematical model already introduced in [14], we interpret its pixels $I_{h,k}$ as the values of an ideal function $f$ at the Chebyshev grid with the same size, namely
$I_{h,k}=f(x_{h}^{N},x_{k}^{M}),\qquad\mbox{with}\quad x_{j}^{n}=\cos\frac{(2j-1)\pi}{2n},\quad j=1,\ldots,n.$
Hence, these pixels are the coefficients of the bivariate VP interpolation polynomial achieved via tensor product, i.e.
$V_{N,M}^{\nu,\mu}f(x,y)=\sum_{h=1}^{N}\sum_{k=1}^{M}I_{h,k}\ \phi_{N,h}^{\nu}(x)\phi_{M,k}^{\mu}(y),\qquad(x,y)\in[-1,1]^{2}$ (24)
where, as usual, we set $\nu=\lfloor\theta N\rfloor$ and $\mu=\lfloor\theta M\rfloor$ with arbitrarily fixed $\theta\in]0,1[$.
Similarly to the univariate case, the VP interpolating polynomial $V_{N,M}^{\nu,\mu}f$ projects the unknown function $f$ onto the approximation space $\mathscr{V}_{N}^{\nu}\times\mathscr{V}_{M}^{\mu}$ and represents a near best polynomial approximation of $f$, as studied in [16].
From the wavelets point of view, the $N\times M$ pixels $I_{h,k}$ constitute the scaling coefficients forming the scaling matrix $I$ to be decomposed and then reconstructed through a standard tensor product approach.
Indeed, supposing for simplicity that both the dimensions $N$ and $M$ are divisible by three, e.g. setting
$N=3N_{1}\qquad\mbox{and}\qquad M=3M_{1}$
the approximation space $\mathscr{V}_{N}^{\nu}\times\mathscr{V}_{M}^{\mu}$ can be decomposed into the orthogonal sum of the lower-degree approximation space and three detail spaces corresponding to the horizontal, vertical, and diagonal directions, i.e.
$\mathscr{V}_{N}^{\nu}\times\mathscr{V}_{M}^{\mu}=\left(\mathscr{V}_{N_{1}}^{\nu}\times\mathscr{V}_{M_{1}}^{\mu}\right)\bigoplus\left(\mathscr{W}_{N_{1}}^{\nu}\times\mathscr{V}_{M}^{\mu}\right)\bigoplus\left(\mathscr{V}_{N}^{\nu}\times\mathscr{W}_{M_{1}}^{\mu}\right)\bigoplus\left(\mathscr{W}_{N_{1}}^{\nu}\times\mathscr{W}_{M_{1}}^{\mu}\right).$
This decomposition of spaces corresponds to a similar decomposition of the VP interpolating polynomial (24) as the sum of one lower-degree approximation and three detail polynomials. In terms of coefficient matrices in the scaling and wavelet bases defined by tensor products, the initial scaling matrix $I$ can therefore be decomposed (and reconstructed) into a scaling matrix at the lower resolution $N_{1}\times M_{1}$ and three detail matrices of size $2N_{1}\times M$ (horizontal details), $N\times 2M_{1}$ (vertical details), and $2N_{1}\times 2M_{1}$ (diagonal details), respectively.
Hence, the decomposition and reconstruction of a matrix $I$ can be carried out using the same one-dimensional algorithms (Algorithms 5 and 6) iteratively applied to the columns and rows of the image data
@@
With the described algorithms and all the above observations in mind, we now proceed to test the VP basis through experiments on two typical data processing tasks: signal denoising and image compression. In the first case, we use univariate VP wavelets, in the second case we deal with bidimensional VP wavelets.
Figure 2: Six test signals: Blocks, Bumps, Heavy Sine, Doppler, Quadchirp, Mishmash.
Figure 3: The results for different values of $\theta=0.1,0.2,\ldots,0.9$ for the six test signals: Blocks, Bumps, Heavy Sine, Doppler, Quadchirp, Mishmash.